# Gradient Methods with Online Scaling Part II. Practical Aspects

### v1

@@ 2 Algorithmic behavior and ingredients of practical OSGM
In this section, we motivate the key ingredients of OSGM-Best that improve the practical convergence of OSGM. We provide insights into each ingredient through either rigorous analysis or illustrative examples. Our analysis focuses on OSGM-H, OSGM with hypergradient feedback, although most of the insights extend naturally to other variants of OSGM. We briefly recap OSGM-H (Algorithm˜1), OSGM using the hypergradient feedback [14]
@@
1 input: $x^{1}$, $P_{1}\in\mathbb{R}^{n\times n}$, online gradient stepsize $\eta>0$
@@ 5 $P_{k+1}=P_{k}-\eta\nabla h_{x^{k}}(P_{k})$
1 input: $x^{1}$, $P_{1}\in\mathcal{P}=\mathbb{R}^{n\times n}$, $\eta>0$
@@ 2.1 Monotone landscape action for hypergradient feedback
The convergence analysis for OSGM-H in Part I [14] always requires a monotone landscape action to ensure at least $f(x^{k+1})\leq f(x^{k})$. In practice, OSGM-H is still observed to converge without enforcing monotonicity, but it exhibits periodic “spikes” in the objective (Figure˜1 (left)). Similar spiky patterns have also been observed in other methods, such as heavy-ball momentum [7] and the Borwein-Barzilai method [11].
@@
The sign of stepsize change $\Delta_{k}$ is determined by both $1-\alpha_{k}$ and $1-\kappa\alpha_{k}$. When $\alpha_{k}>\frac{1}{\kappa}$, $1-\kappa\alpha_{k}<0$ from the second coordinate will try reducing the stepsize, while $1-\alpha_{k}$ from the first coordinate tends to increase the stepsize until $\alpha_{k}\geq 1\gg\alpha^{\star}$. However, these two terms are weighted by quantities that depend on the location of $x^{k}$ on the landscape. In particular, if $x_{2}^{k}\ll\frac{1}{\kappa}x_{1}^{k}$, then $\tfrac{\kappa^{2}(x^{k}_{2})^{2}}{(x^{k}_{1})^{2}+\kappa^{2}(x^{k}_{2})^{2}}\rightarrow 0$, and the information from the second coordinate $1-\kappa\alpha_{k}$ will be disregarded. It can therefore lead to overly large stepsize and divergence. Suppose OSGM-H is initialized such that $x^{k}_{1}\gg\kappa x^{k}_{2}$. Then each spike in Figure˜1 can be decomposed into three phases.
@@
Assume $\kappa$ is large enough to have a constant $a$ such that $\tfrac{2}{\kappa}\ll\tfrac{a}{2}<\tfrac{1}{2}$, say $a=\tfrac{1}{3}$. Note that if the condition in Lemma˜2.1 is always satisfied, OSGM-H takes at most $\lceil\tfrac{4a}{\eta}\rceil$ iterations to increase the stepsize from $\alpha_{1}=0$ to at least $a$.
During the stepsize increasing period, the magnitude $|x^{k}_{2}|$ is at most expanded by a factor $|1-\kappa\alpha_{k}|<\kappa-1$ since $\alpha_{k}<1$. Therefore, we can choose $\delta>0$ small enough such that the iterates $\{(x^{k}_{1},x^{k}_{2})\}_{k\leq K}$ always satisfy the condition in Lemma˜2.1 but $(x^{K+1}_{1},x^{K+1}_{2})$ violates the same condition. Then OSGM-H starts to decrease the stepsize after iteration $K+1$, but the stepsize $\alpha_{K+1}>a\gg\tfrac{2}{\kappa}$ and thus $x_{2}$ continues to expand, causing a spike in the objective. OSGM-H takes $\Theta(\tfrac{1}{\eta\kappa})$ iterations to decrease the stepsize from $a$ to $\tfrac{a}{2}$ since for $\alpha_{k}\leq a$, every iteration can decrease the stepsize by at most $\eta(\kappa a-1)$:
@@ 2.2 Limitations of Classic-HDM
$\begin{array}[]{ll}{{\texttt{Classic-HDM}}}&\texttt{OSGM-H}\\ P_{k+1}=P_{k}-\eta\nabla h_{x^{k}}(P_{k})&x^{k+1}=x^{k}-P_{k}\nabla f(x^{k})\\ x^{k+1}=x^{k}-P_{k+1}\nabla f(x^{k})&P_{k+1}=P_{k}-\eta\nabla h_{x^{k}}(P_{k})\end{array}$
@@
$\begin{array}[]{ll}{{\texttt{Classic-HDM}}}&\texttt{OSGM-H}\\ \alpha_{k+1}=\alpha_{k}+\eta-\alpha_{k}\eta\tfrac{\langle z^{k},\Lambda^{3}z^{k}\rangle}{\langle z^{k},\Lambda^{2}z^{k}\rangle}&z^{k+1}=(I-\alpha_{k}\Lambda)z^{k}\\ z^{k+1}=(I-\alpha_{k+1}\Lambda)z^{k}&\alpha_{k+1}=\alpha_{k}+\eta-\alpha_{k}\eta\tfrac{\langle z^{k},\Lambda^{3}z^{k}\rangle}{\langle z^{k},\Lambda^{2}z^{k}\rangle}\end{array}$
@@
${\mathcal{F}_{{{\texttt{HDM}}}}}(\alpha,\hat{z})\coloneqq\big{(}\alpha^{+}(\hat{z}),\tfrac{(I-\alpha^{+}(\hat{z})\Lambda)\hat{z}}{\|(I-\alpha^{+}(\hat{z})\Lambda)\hat{z}\|}\big{)}\quad\text{and}\quad{\mathcal{F}_{{{\texttt{OSGM}}}}}(\alpha,\hat{z})\coloneqq\big{(}\alpha^{+}(\hat{z}),\tfrac{(I-\alpha\Lambda)\hat{z}}{\|(I-\alpha\Lambda)\hat{z}\|}\big{)}$ (4)
@@
Both Classic-HDM and OSGM-H have additional period-two orbits when the iterates produced fall in a subspace orthogonal to certain eigenvectors of $A$. For example, when $\langle q_{1},z^{1}\rangle=0$, then the orbit can be derived based on $L=\lambda_{2}$. Without loss of generality, we focus on the orbit generated by the iterations that satisfy $|\langle q_{1},z^{k}\rangle|>0,|\langle q_{n},z^{k}\rangle|>0$ for all $k$. In this case, the algorithm iterations always have non-vanishing components along both top and bottom eigen directions.
Although both Classic-HDM and OSGM-H share the same orbits, they exhibit different stability properties around the orbit.
@@ 2.3 Heavy-ball momentum and potential function-based feedback
The heavy-ball feedback (7) depends on two points $x$ and $x^{-}$, which we denote by a state $z=(z_{1},z_{2})=(x,x^{-})$. With slight abuse of notation, we denote $f(z):=f(z_{1})$ and $\nabla f(z):=\nabla f(z_{1})$. We will also slightly abuse the notation of the gradient by directly concatenating the gradients with respect to $P$ and $\beta$: $\nabla h_{z}(P,\beta)=(\nabla_{P}h_{z}(P,\beta),\nabla_{\beta}h_{z}(P,\beta))$. The pushforward operation and the heavy-ball feedback are written with respect to the state $z$ as
@@
the monotone landscape action for OSGM-HB must ensure the potential is nonincreasing:
@@ 2.4 Algorithm design of OSGM-Best
We put together the insights above and propose OSGM-HB, a variant of OSGM that demonstrates competitive performance in practice. OSGM-Best utilizes the lookahead landscape action, tailored to heavy-ball feedback, to suppress regret and thereby accelerate convergence.
Part I [14] introduced the lookahead landscape action, which takes an additional gradient step on the proposed iterate $x^{k+1/2}$ to suppress the regret of online gradient descent on feedback. Since the heavy-ball feedback is defined with respect to the potential function, the additional gradient step should be taken with respect to the potential function $\varphi_{\omega}(\cdot,~z_{2}^{k+1/2})$. As a result, the lookahead landscape action for OSGM-HB is
@@
where $\nabla_{1}\varphi_{\omega}(z)$ denotes the gradient with respect to $z_{1}$ and the stepsize $\frac{1}{L+\omega}$ is the inverse of the smoothness constant $L+\omega$ of the potential function $\varphi_{\omega}(z)$ with respect to $z_{1}$. Monotone Lookahead OSGM-HB determines the next state $z^{k+1}$ by composing the lookahead and monotone landscape actions:
$\varphi_{\omega}(z^{k+1})\leq\min\{\varphi_{\omega}(\mathcal{M}_{\text{look}}(z^{k+1/2})),\varphi_{\omega}(z^{k})\}.$
Algorithm˜3 summarizes Monotone Lookahead OSGM-HB, with monotone landscape action implemented by the null step. The global convergence of Monotone Lookahead OSGM-HB is established in Theorem˜2.3.
1 input Initial state $z^{1}=(z^{1}_{1},z^{1}_{2})$ with $z_{1}^{1}=z_{2}^{1}$, initial stepsize $P_{1}\in\mathbb{R}^{n\times n}$ and momentum parameter $\beta_{1}\in\mathbb{R}$, online gradient descent stepsizes $\eta_{P},\eta_{\beta}>0$
@@
5   Choose $z^{k+1}$ satisfying $\varphi_{\omega}(z^{k+1})\leq\min\{\varphi_{\omega}(z_{\text{look}}^{k+1}),\varphi_{\omega}(z^{k})\}$ $\triangleright$  Monotone action by null step
@@ 3 OSGM for smooth nonconvex optimization
OSGM can be extended to smooth nonconvex optimization problems, but requires modifying the feedback. The analysis of OSGM in Part I [14] and Section˜2 assumes a convex objective $f$, which guarantees convexity of feedback and enables hypergradient [14, Theorem 4.2] and heavy-ball (Theorem˜2.2) progress reduction.
One possible approach to handle nonconvex $f$ is to apply the proximal point method [4] with a sufficiently large regularizer $\gamma>L$:
@@ 3.3 Algorithm design and analysis
$\ell_{x}(P)\coloneqq h_{x}^{\lambda}(P)\quad\text{Monotone Lookahead landscape: }f(x^{k+1})\leq\min\{f(x^{k+1/2}-\tfrac{1}{L}\nabla f(x^{k+1/2})),f(x^{k})\}\quad\mathcal{A}\coloneqq{\texttt{OGD}}.$
@@
1 input initial point $x^{1}$, initial stepsize $P_{1}\in\mathcal{P}$, online gradient stepsize $\eta>0$
@@ 5.3 OSGM for accelerated gradient descent
1 input initial point $x^{1}=z^{1}$
@@ 4 $x^{k+1}=y^{k}-\tfrac{1}{L}\nabla f(y^{k})$
Algorithm 5 Accelerated gradient descent
1 input initial point $x^{1}=z^{1}$
@@ Appendix A Proof of results in Section˜2
The proof of Theorem˜2.1 requires several lemmas. The proof of each lemma is deferred to the the end of the section.
@@
Let $I_{1,n}\coloneqq[e_{1},~e_{n}]\in\mathbb{R}^{n\times 2}$ denote the first and last column of an $n\times n$ identity matrix.
Using (35) and (36), we have
@@
We can simultaneously permute the rows and columns so that all the four Jacobians are similar to a block diagonal matrix consists of n2n-2 individual diagonal entries $1-\alpha^{\star}\lambda_{2},\ldots,1-\alpha^{\star}\lambda_{n-1}$ and one $3\times 3$ dense matrix, denoted by $J_{{{\texttt{HDM}}}}^{i}$ and $J_{{\texttt{OSGM}}}^{i}$ for $i=1,2$. After careful but tedious calculations, the four $3\times 3$ matices match the their definition in the lemma.
@@ Appendix C Proof of results in Section 5
1 input initial point $x^{1}$, initial stepsize $P_{1}\in\mathbb{R}^{n\times n}$, online proximal regularizer $\eta>0$
@@ 2. A2:
1 input initial point $x^{1}$, initial stepsize $P_{1}\in\mathbb{R}^{n\times n}$, online proximal regularizer $\eta>0$

### v2

@@ 2 Algorithmic behavior and ingredients of practical OSGM
In this section, we motivate the key ingredients of OSGM-Best that improve the practical convergence of OSGM. We provide insights into each ingredient through either rigorous analysis or illustrative examples. Our analysis focuses on OSGM-H (Algorithm˜1), OSGM using hypergradient feedback, although most of the insights extend naturally to other variants of OSGM. The hypergradient feedback [14] takes the form
@@
1input: $x^{1}$, $P_{1}\in\mathbb{R}^{n\times n}$, online gradient stepsize $\eta>0$
@@ 5 $P_{k+1}=P_{k}-\eta\nabla h_{x^{k}}(P_{k})$
1input: $x^{1}$, $P_{1}\in\mathcal{P}=\mathbb{R}^{n\times n}$, $\eta>0$
@@ 2.1 Monotone landscape action for hypergradient feedback
The convergence analysis for OSGM-H in Part I [14] always requires a monotone landscape action to ensure at least $f(x^{k+1})\leq f(x^{k})$. In practice, OSGM-H is still observed to converge without enforcing monotonicity, but it exhibits periodic “spikes” in the objective (Figure˜1 (left)). Similar spiky patterns have also been observed in other methods, such as heavy-ball momentum [7] and the Barzilai-Borwein method [11].
@@
The sign of stepsize change $\Delta_{k}$ is determined by both $1-\alpha_{k}$ and $1-\kappa\alpha_{k}$. When $\alpha_{k}>\frac{1}{\kappa}$, the value $1-\kappa\alpha_{k}<0$ from the second coordinate will try reducing the stepsize, while $1-\alpha_{k}$ from the first coordinate tends to increase the stepsize until $\alpha_{k}\geq 1\gg\alpha^{\star}$. However, these two terms are weighted by quantities that depend on the location of $x^{k}$ on the landscape. In particular, if $x^{k}_{1}\gg\kappa x^{k}_{2}$, then $\tfrac{\kappa^{2}(x^{k}_{2})^{2}}{(x^{k}_{1})^{2}+\kappa^{2}(x^{k}_{2})^{2}}$ is close to 0, and the value $1-\kappa\alpha_{k}$ from the second coordinate will be disregarded. It can therefore lead to overly large stepsize and divergence. Suppose OSGM-H is initialized such that $x^{k}_{1}\gg\kappa x^{k}_{2}$. Then each spike in Figure˜1 can be decomposed into three phases.
@@
Assume $\kappa$ is large enough to have a constant $a$ such that $\tfrac{2}{\kappa}\ll\tfrac{a}{2}<\tfrac{1}{2}$, say $a=\tfrac{1}{3}$. Note that if the condition in Lemma˜2.1 is always satisfied, OSGM-H takes at most $\lceil\tfrac{4a}{\eta}\rceil$ iterations to increase the stepsize from $\alpha_{1}=0$ to at least $a$. During the stepsize increasing period, the magnitude $|x^{k}_{2}|$ is at most expanded by a factor $|1-\kappa\alpha_{k}|<\kappa-1$ since $\alpha_{k}<1$. Therefore, we can choose $\delta>0$ small enough such that the iterates $\{(x^{k}_{1},x^{k}_{2})\}_{k\leq K}$ always satisfy the condition in Lemma˜2.1 but $(x^{K+1}_{1},x^{K+1}_{2})$ violates the same condition. Then OSGM-H starts to decrease the stepsize after iteration $K+1$, but the stepsize $\alpha_{K+1}>a\gg\tfrac{2}{\kappa}$ and thus $x_{2}$ continues to expand, causing a spike in the objective. OSGM-H takes $\Theta(\tfrac{1}{\eta\kappa})$ iterations to decrease the stepsize from $a$ to $\tfrac{a}{2}$ since for $\alpha_{k}\leq a$, every iteration can decrease the stepsize by at most $\eta(\kappa a-1)$:
@@ 2.2 Limitations of Classic-HDM
$\begin{array}[]{ll}{{\texttt{Classic-HDM}}}&\texttt{OSGM-H}\\ P_{k+1}=P_{k}-\eta\nabla h_{x^{k}}(P_{k})&x^{k+1}=x^{k}-P_{k}\nabla f(x^{k})\\ x^{k+1}=x^{k}-P_{k+1}\nabla f(x^{k})\qquad&P_{k+1}=P_{k}-\eta\nabla h_{x^{k}}(P_{k})\end{array}$
@@
$\begin{array}[]{ll}{{\texttt{Classic-HDM}}}&\texttt{OSGM-H}\\ \alpha_{k+1}=\alpha_{k}+\eta-\alpha_{k}\eta\tfrac{\langle z^{k},\Lambda^{3}z^{k}\rangle}{\langle z^{k},\Lambda^{2}z^{k}\rangle}\qquad&z^{k+1}=(I-\alpha_{k}\Lambda)z^{k}\\ z^{k+1}=(I-\alpha_{k+1}\Lambda)z^{k}\qquad&\alpha_{k+1}=\alpha_{k}+\eta-\alpha_{k}\eta\tfrac{\langle z^{k},\Lambda^{3}z^{k}\rangle}{\langle z^{k},\Lambda^{2}z^{k}\rangle}\end{array}$
@@
[…]
@@
Both Classic-HDM and OSGM-H have additional period-two orbits when the iterates fall in a subspace orthogonal to certain eigenvectors of $A$. For example, when $\langle q_{1},z^{1}\rangle=0$, then the orbit can be derived based on $L=\lambda_{2}$ (i.e. the maximum eigenvalue restricted on subspace). Without loss of generality, we focus on the orbit generated by the iterations that satisfy $|\langle q_{1},z^{k}\rangle|>0,|\langle q_{n},z^{k}\rangle|>0$ for all $k$. In this case, the algorithm iterations always have non-vanishing components along both top and bottom eigen directions.
Although both Classic-HDM and OSGM-H share the same orbits, Theorem˜2.1 shows that they exhibit different stability properties around the orbit. The proof of Theorem˜2.1 is deferred to appendix due to its length and technicality.
@@ 2.3 Heavy-ball momentum and potential function-based feedback
The heavy-ball feedback (7) depends on two points $x$ and $x^{-}$, which we denote by a state $z=(z_{1},z_{2})=(x,x^{-})$. With slight abuse of notation, we denote $f(z):=f(z_{1})$ and $\nabla f(z):=\nabla f(z_{1})$. We will also slightly abuse the notation of the gradient by directly concatenating the gradients with respect to $P$ and $\beta$: $\nabla h_{z}(P,\beta)=(\nabla_{P}h_{z}(P,\beta),\nabla_{\beta}h_{z}(P,\beta))$. The pushforward operation and the heavy-ball feedback are written with respect to state $z$ as
@@
the monotone landscape action for heavy-ball feedback must ensure the potential is nonincreasing:
@@ 2.4 Algorithm design of OSGM-Best
We put together the insights above and propose OSGM-Best, a variant of OSGM that demonstrates competitive performance in practice. OSGM-Best utilizes the lookahead landscape action, tailored to heavy-ball feedback, to suppress regret and thereby accelerate convergence.
Part I [14] introduced the lookahead landscape action, which takes an additional gradient step on the proposed iterate $x^{k+1/2}$ to suppress the regret of online gradient descent on feedback. Since the heavy-ball feedback is defined with respect to the potential function, the additional gradient step should be taken with respect to the potential function $\varphi_{\omega}(\cdot,~z_{2}^{k+1/2})$. As a result, the lookahead landscape action for heavy-ball feedback is
@@
where $\nabla_{1}\varphi_{\omega}(z)$ denotes the gradient with respect to $z_{1}$ and the stepsize $\frac{1}{L+\omega}$ is the inverse of the smoothness constant $L+\omega$ of the potential function $\varphi_{\omega}(z)$ with respect to $z_{1}$. OSGM-Best determines the next state $z^{k+1}$ by composing the lookahead and monotone landscape actions:
Algorithm˜3 summarizes OSGM-Best with monotone landscape action implemented by the null step. We refer to Algorithm˜3 as OSGM-Best since it is empirically superior to other OSGM variants. The global convergence of OSGM is established in Theorem˜2.3.
1input Initial state $z^{1}=(z^{1}_{1},z^{1}_{2})$ with $z_{1}^{1}=z_{2}^{1}$, initial stepsize $P_{1}\in\mathbb{R}^{n\times n}$ and momentum parameter $\beta_{1}\in\mathbb{R}$, online gradient descent stepsizes $\eta_{P},\eta_{\beta}>0$
@@
5   Choose $z^{k+1}$ s.t. $\varphi_{\omega}(z^{k+1})\leq\min\{\varphi_{\omega}(z_{\text{look}}^{k+1}),\varphi_{\omega}(z^{k})\}$ $\triangleright$  Monotone action by null step
@@ 3 OSGM for smooth nonconvex optimization
OSGM can be extended to smooth nonconvex optimization problems, but requires modifying the feedback. The analysis of OSGM in Part I [14] and Section˜2 assumes a convex objective $f$, which guarantees convexity of feedback and enables hypergradient [14, Theorem 4.2] and heavy-ball (Theorem˜2.2) progress reduction. One possible approach to handle nonconvex $f$ is to apply the proximal point method [4] with a sufficiently large regularizer $\gamma>L$:
@@ 3.3 Algorithm design and analysis
$\ell_{x}(P)\coloneqq h_{x}^{\lambda}(P),\quad\text{Monotone Lookahead landscape: }f(x^{k+1})\leq\min\{f(x^{k+1/2}-\tfrac{1}{L}\nabla f(x^{k+1/2})),f(x^{k})\}\quad\mathcal{A}\coloneqq{\texttt{OGD}}.$
@@
1input initial point $x^{1}$, initial stepsize $P_{1}\in\mathcal{P}$, online gradient stepsize $\eta>0$
@@ 5.3 OSGM for accelerated gradient descent
1input initial point $x^{1}=z^{1}$
@@ 4 $x^{k+1}=y^{k}-\tfrac{1}{L}\nabla f(y^{k})$
Algorithm 5 Accelerated GD
1input initial point $x^{1}=z^{1}$
@@ Appendix A Proof of results in Section˜2
The proof of Theorem˜2.1 requires several lemmas. The proof of each lemma is deferred to the end of the section.
@@
Let $I_{1,n}\coloneqq[e_{1},~e_{n}]\in\mathbb{R}^{n\times 2}$ denote the first and last column of an $n\times n$ identity matrix. Using (36) and (37), we have
@@
We can simultaneously permute the rows and columns so that all the four Jacobians are similar to a block diagonal matrix that consists of n2n-2 individual diagonal entries $1-\alpha^{\star}\lambda_{2},\ldots,1-\alpha^{\star}\lambda_{n-1}$ and one $3\times 3$ dense matrix, denoted by $J_{{{\texttt{HDM}}}}^{i}$ and $J_{{\texttt{OSGM}}}^{i}$ for $i=1,2$. After careful but tedious calculations, the four $3\times 3$ matrices match their definition in the lemma.
@@ Appendix C Proof of results in Section 5
1input initial point $x^{1}$, initial stepsize $P_{1}\in\mathbb{R}^{n\times n}$, online proximal regularizer $\eta>0$
@@ 2. A2:
1input initial point $x^{1}$, initial stepsize $P_{1}\in\mathbb{R}^{n\times n}$, online proximal regularizer $\eta>0$